OpenStudy (sodapop):

The function f(x) = 16(5)x represents the growth of a bee population every year in a remote swamp. Jennifer wants to manipulate the formula to an equivalent form that calculates two times a year, not just once a year. Which function is correct for Jennifer's purpose, and what is the new growth rate?

OpenStudy (sodapop):

@hero @Jim766

OpenStudy (sodapop):

f(x)=8(5)^x; growth rate 500% f(x)=8(5)2^x; growth rate 8% f(x)=8(2.24)^x; growth rate 8% f(x)=8(2.24)^2x; growth rate 224%

OpenStudy (jim766):

5^x=t^2x, 5^x=(t2)^x t^2=5,t=5√=2.24 new rate=2.24 i think it can be 4th but why 8 in place of 16,i don't know.

OpenStudy (sodapop):

Alright so that gets rid of answers A and B. Now how do we find the growth rate?

OpenStudy (jim766):

2.24 * 100 = 224%

OpenStudy (sodapop):

Alright sweet, thank you.

hero (hero):

I still assert that the original f(x) is NOT equivalent to any of the answer choices. You can simply graph each equation and see for yourself.

OpenStudy (sodapop):

Thank you